# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 3 Observability Problems
By applying Carleman estimate (1.6) with
we have that for $\lambda\geq C(T+T^{2})$,
First, it is easy to see that
We also have that for a large $\lambda\geq CT^{2}$,
From Corollary 3.1, we first deduce the following unique continuation property for the system (3.1).
Proposition 3.1.
[…]
$(z,Z,\widehat{Z})=(0,0,0)\,\,\,\textnormal{in}\,\,\,Q_{0}\times Q\times\Sigma,\;\,\textnormal{a.s.}\Longrightarrow\;\,(z_{T},z_{\Gamma,T})=(0,0),\,\,\,\textnormal{in}\,\,\,G\times\Gamma,\;\,\textnormal{a.s.}$ (3.7)
Secondly, from Corollary 3.1, we also derive the following observability inequality.
Proposition 3.2.
[…]
$\begin{aligned} &\,\|z(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(G))}+\|z_{\Gamma}(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\Gamma))}\\ &\leq e^{CK}\,\bigg[\mathbb{E}\iint_{Q_{0}}z^{2}dxdt+\mathbb{E}\iint_{Q}Z^{2}dxdt+\mathbb{E}\iint_{\Sigma}\widehat{Z}\,^{2}d\sigma dt\bigg],\end{aligned}$ (3.8)
where $K$ has the following form:
Proof.
From the Carleman inequality (3.3), we deduce
for large $\lambda\geq\lambda_{1}(T)\equiv\lambda_{0}\left[T+T^{2}(1+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2})\right].$ Subsequently, notice that the functions
$t\mapsto\min_{x\in\overline{G}}\,[\theta^{2}(t,x)\varphi^{3}(t,x)]$
reach their minimum in $[T/4,3T/4]$ at $t=T/4$ and
$t\mapsto\max_{x\in\overline{G}}\,[\theta^{2}(t,x)\varphi^{3}(t,x)]$
On the other hand, let $t\in(0,T)$, computing $d_{\tau}\|(z,z_{\Gamma})\|^{2}_{\mathbb{L}^{2}}$ by using the Itô’s formula. Then integrating the equality w.r.t $\tau\in(0,t)$ and taking the expectation on both sides, we conclude that
@@
$\begin{aligned} &\,\mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma-\left(\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\right)\\ &=-2\mathbb{E}\int_{0}^{t}\int_{G}\mathcal{A}\nabla z\cdot\nabla zdxd\tau+2\mathbb{E}\int_{0}^{t}\int_{G}a_{1}z^{2}dxd\tau+2\mathbb{E}\int_{0}^{t}\int_{G}zB_{1}\cdot\nabla z\,dxd\tau\\ &\quad\,-\mathbb{E}\int_{0}^{t}\int_{G}Z^{2}dxd\tau-2\mathbb{E}\int_{0}^{t}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma d\tau+2\mathbb{E}\int_{0}^{t}\int_{\Gamma}a_{2}z_{\Gamma}^{2}d\sigma d\tau\\ &\quad\,+2\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}B_{2}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma d\tau-\mathbb{E}\int_{0}^{t}\int_{\Gamma}\widehat{Z}^{2}d\sigma d\tau.\end{aligned}$
$\begin{aligned} &\,\mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma-\left(\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\right)\\ &\leq-2\beta_{0}\mathbb{E}\int_{0}^{t}\int_{G}|\nabla z|^{2}dxd\tau+2\|a_{1}\|_{\infty}\mathbb{E}\int_{0}^{t}\int_{G}z^{2}dxd\tau+\varepsilon\mathbb{E}\int_{0}^{t}\int_{G}|\nabla z|^{2}\,dxd\tau\\ &\quad\,+\frac{\|B_{1}\|_{\infty}^{2}}{\varepsilon}\mathbb{E}\int_{0}^{t}\int_{G}z^{2}\,dxd\tau-\mathbb{E}\int_{0}^{t}\int_{G}Z^{2}dxd\tau-2\beta_{0}\mathbb{E}\int_{0}^{t}\int_{\Gamma}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma d\tau\\ &\quad\,+2\|a_{2}\|_{\infty}\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}^{2}d\sigma d\tau+\varepsilon\mathbb{E}\int_{0}^{t}\int_{\Gamma}|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma d\tau+\frac{\|B_{2}\|_{\infty}^{2}}{\varepsilon}\mathbb{E}\int_{0}^{t}\int_{\Gamma}z^{2}_{\Gamma}\,d\sigma d\tau\\ &\quad\,-\mathbb{E}\int_{0}^{t}\int_{\Gamma}\widehat{Z}^{2}d\sigma d\tau.\end{aligned}$ (3.11)
Taking a small $\varepsilon$ in (3.11), we conclude that
$\begin{aligned} \mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma&\leq\,\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\\ &\hskip 11.38092pt+C\textbf{K}_{2}\left(\mathbb{E}\int_{0}^{t}\int_{G}z^{2}dxd\tau+\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}^{2}d\sigma d\tau\right),\end{aligned}$
4 Controllability Results: Proof of Theorems 1.2 and 1.3
This section is devoted to proving the null and approximate controllability results for the forward equation (1.7), as stated in Theorems 1.2 and 1.3, respectively. To proceed with the proofs, we first recall the following well-known result from functional analysis (see, e.g., [30] and [36, Theorem IV.2.2]).
Lemma 4.1.
if and only if
Moreover, the constant $\mathcal{C}$ in (4.1) and (4.2) is the same.
Let us first prove the approximate controllability of the equation (1.7).
Proof of Theorem 1.2.
Define the following controllability operator
$\mathcal{L}_{T}(u,v_{1},v_{2})=(y(T,\cdot),y_{\Gamma}(T,\cdot)),$
where $(y,y_{\Gamma})$ is the solution of (1.7) with initial data $(y_{0},y_{\Gamma,0})=(0,0)$. It is easy to see that the equation (1.7) is approximately controllable at time $T$ if and only if $\mathcal{R}(\mathcal{L}_{T})$ is dense in $L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, equivalently the adjoint operator $\mathcal{L}_{T}^{*}$ is injective. Let us first compute $\mathcal{L}_{T}^{*}$, for this, multiplying the equations (3.1) and (1.7) using Itô’s formula, we find the following duality relation
It follows that for all $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, the adjoint operator $\mathcal{L}_{T}^{*}$ is given by
[…]
Hence, (1.7) is approximately controllable at time $T$ if and only if the property
$(z,Z,\widehat{Z})=(0,0,0)\,\,\,\textnormal{in}\,\,\,Q_{0}\times Q\times\Sigma,\;\,\textnormal{a.s.}\;\;\Longrightarrow\;\,(z_{T},z_{\Gamma,T})=(0,0),\,\,\,\textnormal{in}\,\,\,G\times\Gamma,\;\,\textnormal{a.s.},$ (4.5)
holds. Since, the property (4.5) is achieved in Proposition 3.1, therefore the equation (1.7) is approximately controllable at time $T$. This completes the proof of Theorem 1.2. ∎
Let us now show the null controllability of (1.7).
Proof of Theorem 1.3.
Let us first define the following linear bounded operator
where $(y,y_{\Gamma})$ is the solution of (1.7) with controls $(u,v_{1},v_{2})\equiv(0,0,0)$. Also, we consider the following controllability operator
$\mathcal{L}_{T}(u,v_{1},v_{2})=(y(T,\cdot),y_{\Gamma}(T,\cdot)),$
where $(y,y_{\Gamma})$ is the solution of (1.7) with initial data $(y_{0},y_{\Gamma,0})\equiv(0,0)$.
Recalling (4.3) for the duality relation between the equation (1.7) and its adjoint equation (3.1), it is easy to see that the adjoint operators of $\mathcal{M}_{T}$ and $\mathcal{L}_{T}$ are respectively defined by
[…]
where
Let $(y_{0},y_{\Gamma,0})\in{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}$ be the initial state of (1.7). From Lemma 4.1 and the inequality (4.7), we deduce that $\mathcal{R}(\mathcal{M}_{T})\subset\mathcal{R}(\mathcal{L}_{T})$, and hence there exists controls $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$ such that
On the other hand, it is obvious that (1.7) is null controllable at time $T$ if and only if $\mathcal{R}(\mathcal{M}_{T})\subset\mathcal{R}(\mathcal{L}_{T})$. Therefore, the equation (1.7) is null controllable at time $T$, and from (4.8), the desired estimate (1.8) holds as well. This concludes the proof of Theorem 1.3. ∎
Appendix A Proof of Lemma 2.1
In this appendix, we follow some computations outlined in the proof of [6, Theorem 4.1] and [3, Lemma 3.2]. However, in our case, we must carefully handle the dependence of Carleman’s parameters on the final control time $T$, as well as work with the more general diffusion coefficients $\mathcal{A}$ and $\mathcal{A}_{\Gamma}$, instead of Laplacian operators.
First, we recall that the diffusion matrix operator $\mathcal{A}=(a^{jk})_{j,k=1,2,\cdots,N}$ satisfies the following assumptions:
  1. 1.
    $a^{jk}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{2,\infty}(\overline{G})))$, and $a^{jk}=a^{kj}$ for all $j,k=1,2,\cdots,N$.
  2. 2.
    There exists a constant $\beta_{0}>0$ such that
    $\sum_{j,k=1}^{N}a^{jk}(\omega,t,x)\xi_{j}\xi_{k}\geq\beta_{0}|\xi|^{2},\quad\textnormal{for all}\;\;(\omega,t,x,\xi)\equiv(\omega,t,x,\xi_{1},\cdots,\xi_{N})\in\Omega\times(0,T)\times\overline{G}\times\mathbb{R}^{N}.$
Define a function $\ell\in C^{1,3}((0,T)\times\overline{G})$, and two auxiliary functions $\Psi\in C^{1,2}((0,T)\times\overline{G})$ and $\Phi\in C^{1,0}(\Sigma)$. For the sake of simplicity in notation, we let $y_{x_{j}}=\frac{\partial y}{\partial x_{j}}$ denote the partial derivative of the function $y$ with respect to the $j$-th component $x_{j}$ of a generic point $x=(x_{1},\cdots,x_{N})\in\mathbb{R}^{N}$. Similarly, we use $y_{x_{j}x_{k}}$ to denote the second partial derivative of $y$ with respect to the variables $x_{j}$ and $x_{k}$, and so on. Put
$\begin{cases}A=\displaystyle\sum_{j,k=1}^{N}(a^{jk}\ell_{x_{j}}\ell_{x_{k}}-a^{jk}_{x_{k}}\ell_{x_{j}}-a^{jk}\ell_{x_{j}x_{k}})-\Psi-\ell_{t},\\ B=2\Bigg[A\Psi+\displaystyle\sum_{j,k=1}^{N}(Aa^{jk}\ell_{x_{j}})_{x_{k}}\Bigg]-A_{t}+\sum_{j,k=1}^{N}(a^{jk}\Psi_{x_{k}})_{x_{j}},\\ c^{jk}=\displaystyle\sum_{j^{\prime},k^{\prime}=1}^{N}\Big[2a^{jk^{\prime}}(a^{j^{\prime}k}\ell_{x_{j^{\prime}}})_{x_{k^{\prime}}}-(a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}})_{x_{k^{\prime}}}\Big]+\frac{a^{jk}_{t}}{2}-\Psi a^{jk},\quad j,k=1,2,\cdots,N,\\ \widetilde{A}=\Phi-\ell_{t},\\ \widetilde{B}=-2\widetilde{A}\Phi-\widetilde{A}_{t}.\end{cases}$ (A.1)
From [27, Theorem 9.26], we have the following fundamental weighted identity for the stochastic parabolic-like operator $dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt$.
Theorem A.1.
Let $z$ be an $H^{2}(G)$-valued Itô process. Set $\theta=e^{\ell}$ and $h=\theta z$. Then, for any $t\in[0,T]$ and a.e. $(x,\omega)\in\overline{G}\times\Omega$,
 $\begin{aligned} &\,2\theta\left[\nabla\cdot(\mathcal{A}\nabla h)+Ah\right]\left[dz+\nabla\cdot(\mathcal{A}\nabla z)dt\right]-2\nabla\cdot(\mathcal{A}\nabla h\,dh)\\ &+2\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dt\\ &=2\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dt+Bh^{2}dt+d\Bigg(-\sum_{j,k=1}^{N}a^{jk}h_{x_{j}}h_{x_{k}}+Ah^{2}\Bigg)\\ &\;\;\;\,+2\big[\nabla\cdot(\mathcal{A}\nabla h)+Ah\big]^{2}dt+\theta^{2}\sum_{j,k=1}^{N}a^{jk}(dz_{xj}+\ell_{x_{j}}dz)(dz_{x_{k}}+\ell_{x_{k}}dz)\\ &\;\;\;\,-\theta^{2}A(dz)^{2}.\end{aligned}$ (A.2)
We also have the following fundamental weighted identity for the stochastic parabolic-like operator $dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt$.
Theorem A.2.
Let $z_{\Gamma}$ be an $H^{2}(\Gamma)$-valued Itô process. Set $\theta=e^{\ell}$ and $h_{\Gamma}=\theta z_{\Gamma}$ with $z_{\Gamma}=z|_{\Gamma}$. Then, for any $t\in[0,T]$ and a.e. $(x,\omega)\in\Gamma\times\Omega$,
 $\begin{aligned} &\,2\theta\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt\right]-2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\,dh_{\Gamma})\\ &=\widetilde{B}h_{\Gamma}^{2}dt+d(-\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}+\widetilde{A}h_{\Gamma}^{2})+2[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}]^{2}dt\\ &\quad+\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}dt+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}-\theta^{2}\widetilde{A}|dz_{\Gamma}|^{2}-2\nabla_{\Gamma}\cdot(\Phi h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt\\ &\quad+2\Phi\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt+2h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi dt-2\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})dt-2\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zdt.\end{aligned}$ (A.3)
Proof.
We compute the first term on the left-hand side of (A.3). First, by the definition of $\theta$, we have
 $\begin{aligned} &\,2\theta\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+\widetilde{A}h_{\Gamma}dt+dh_{\Gamma}-\Phi h_{\Gamma}dt-\theta\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2I(I_{1}+I_{2}+I_{3}+I_{4}),\end{aligned}$ (A.4)
@@
$2II_{1}=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]^{2}dt.$ (A.5)
By Itô’s formula, we derive that
@@
which implies that
@@
$\begin{aligned} 2II_{2}=&2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}dh_{\Gamma})-d(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma})+\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}dt\\ &+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}+d(\widetilde{A}h_{\Gamma}^{2})-\widetilde{A}_{t}h_{\Gamma}^{2}dt-\theta^{2}\widetilde{A}|dz_{\Gamma}|^{2}.\end{aligned}$ (A.6)
Similarly, we also obtain that
$2II_{3}=-2\nabla_{\Gamma}\cdot(\Phi h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+2\Phi\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt+2h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi-2\widetilde{A}\Phi h_{\Gamma}^{2}dt.$ (A.7)
Finally, we have that
$2II_{4}=-2\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})dt-2\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zdt.$ (A.8)
Combining (A.4), (A.5), (A.6), (A.7), and (A.8), we deduce the desired identity (A.3). ∎
In what follows, we choose $\Psi=-2\sum_{j,k=1}^{N}a^{jk}\ell_{x_{j}x_{k}}$, $\Phi=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}(\mathcal{A}\nu\cdot\nu)|\nabla\ell|$ and $\ell$ is the function defined in (1.4). For a positive integer $n$, we denote by $O(\mu^{n})$ a function of order $\mu^{n}$ for large $\mu$, and similarly, we use the notation $O(e^{\mu|\psi|_{\infty}})$. By (1.4) and (1.5), it is easy to see that for any parameters $\lambda,\mu\geq 1$, we have for any $j,k=1,2,\cdots,N$,
$\ell_{t}=\lambda\alpha_{t},\quad\ell_{x_{j}}=\lambda\mu\varphi\psi_{x_{j}},\quad\ell_{x_{j}x_{k}}=\lambda\mu^{2}\varphi\psi_{x_{j}}\psi_{x_{k}}+\lambda\mu\varphi\psi_{x_{j}x_{k}},$ (A.9)
and also
@@
$\begin{aligned} &\alpha_{t}=T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),\quad\varphi_{t}=T\varphi^{2}O(1),\\ &\alpha_{tt}=T^{2}\varphi^{3}O(e^{2\mu|\psi|_{\infty}}),\quad\varphi_{tt}=T^{2}\varphi^{3}O(1).\end{aligned}$ (A.10)
We have the following estimates for $A$, $B$, and $c^{jk}$ defined in (A.1) (see [3, Theorem A.2]).
Lemma A.1.
@@
There exists a constant $C=C(G,G_{0},\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that for sufficiently large $\mu\geq C$, and $\lambda\geq CT^{2}$, we have that
$A$ $=\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),$
$B\geq$  $\,2\beta_{0}^{2}\lambda^{3}\mu^{4}\varphi^{3}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu|\psi|_{\infty}}),$
@@
$\sum_{j,k=1}^{N}c^{jk}\xi_{j}\xi_{k}\geq$ $\,\left[\beta_{0}^{2}\lambda\mu^{2}\varphi|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right]|\xi|^{2},\quad\forall\xi=(\xi_{1},\cdots,\xi_{N})\in\mathbb{R}^{N},$
@@
for a.e. $(x,\omega)\in\overline{G\setminus G_{1}}\times\Omega$ and $t\in[0,T]$.
@@
Similarly to the above lemma, for the quantities $\widetilde{A}$ and $\widetilde{B}$, we have the following result.
Lemma A.2.
@@
$\widetilde{A}=-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),$
$\widetilde{B}=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}),$

### v2

@@ 3 Intermediate Carleman Estimate without Divergence Terms
We first compute the leading term on the left-hand side of (3.4). By the definition of $\theta$, we have
@@
 $\begin{aligned} &\,2\theta\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+\widetilde{A}h_{\Gamma}dt+dh_{\Gamma}-\Phi h_{\Gamma}dt-\theta\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2I(I_{1}+I_{2}+I_{3}+I_{4}),\end{aligned}$ (3.5)
@@
$2II_{1}=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]^{2}dt.$ (3.6)
By Itô’s formula, we obtain
@@
which can be equivalently written as
@@
$\begin{aligned} 2II_{2}=&2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}dh_{\Gamma})-d(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma})+(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt\\ &+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}+d(\widetilde{A}h_{\Gamma}^{2})-\widetilde{A}_{t}h_{\Gamma}^{2}dt-\theta^{2}\widetilde{A}|dz_{\Gamma}|^{2}.\end{aligned}$ (3.7)
Similarly, we have
$2II_{3}=-2\nabla_{\Gamma}\cdot(\Phi h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+2\Phi\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt+2h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi-2\widetilde{A}\Phi h_{\Gamma}^{2}dt.$ (3.8)
Finally, for the boundary coupling term, we have
$2II_{4}=-2\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})dt-2\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zdt.$ (3.9)
Combining (3.5), (3.6), (3.7), (3.8), and (3.9), we deduce the desired identity (3.4). ∎
In what follows, we choose the auxiliary functions
$\Psi=-2\sum_{j,k=1}^{N}a^{jk}\ell_{x_{j}x_{k}},\qquad\Phi=\frac{2}{\beta_{0}}\,M\,(\mathcal{A}\nu\cdot\nu)\,|\nabla\ell|,$
where the function $\ell$ is chosen to be the one defined in (1.6). Moreover, in Lemma 1.1, the set $G_{1}$ is taken to be any fixed nonempty open subset satisfying $G_{1}\Subset\mathcal{B}$, where $\mathcal{B}$ is any nonempty open subset of $G$. For a positive integer $n$, we denote by $O(\mu^{n})$ a function of order $\mu^{n}$ for sufficiently large $\mu$. Similarly, the notation $O\!\left(e^{\mu\|\psi\|_{\infty}}\right)$ stands for a generic quantity of order $e^{\mu\|\psi\|_{\infty}}$. By (1.6) and (1.7), it is straightforward to verify that for any parameters $\lambda,\mu\geq 1$ and for all $j,k=1,2,\cdots,N$, one has
$\ell_{t}=\lambda\alpha_{t},\qquad\ell_{x_{j}}=\lambda\mu\varphi\psi_{x_{j}},\qquad\ell_{x_{j}x_{k}}=\lambda\mu^{2}\varphi\psi_{x_{j}}\psi_{x_{k}}+\lambda\mu\varphi\psi_{x_{j}x_{k}}.$ (3.10)
Moreover, the following estimates hold:
@@
$\begin{aligned} &\alpha_{t}=T\varphi^{2}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right),\qquad\varphi_{t}=T\varphi^{2}\,O(1),\\[1.99997pt] &\alpha_{tt}=T^{2}\varphi^{3}\,O\!\left(e^{2\mu\|\psi\|_{\infty}}\right),\qquad\varphi_{tt}=T^{2}\varphi^{3}\,O(1).\end{aligned}$ (3.11)
We next give some useful estimates for the quantities $A$, $B$, and $c^{jk}$ defined in (3.2) (see [4, Theorem A.2]).
Lemma 3.1.
@@
There exists a constant $C=C(G,\mathcal{B},\beta_{0},M,M_{\Gamma})>0$ such that, for all $\mu\geq C$ and $\lambda\geq CT^{2}$, the following estimates hold:
$A$ $=\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu\|\psi\|_{\infty}}),$
$B\geq$  $\,2\beta_{0}^{2}\lambda^{3}\mu^{4}\varphi^{3}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu\|\psi\|_{\infty}}),$
@@
$\sum_{j,k=1}^{N}c^{jk}\eta_{j}\eta_{k}\geq$ $\,\left[\beta_{0}^{2}\lambda\mu^{2}\varphi|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right]|\eta|^{2},\quad\forall\eta=(\eta_{1},\cdots,\eta_{N})\in\mathbb{R}^{N},$
@@
for almost every $(x,\omega)\in\overline{G\setminus G_{1}}\times\Omega$ (where $G_{1}$ is defined in Lemma 1.1) and for all $t\in[0,T]$.
@@
Similarly, for the boundary quantities $\widetilde{A}$ and $\widetilde{B}$ defined in (3.2), we have the following estimates.
Lemma 3.2.
@@
$\widetilde{A}=-\frac{2}{\beta_{0}}M\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu\|\psi\|_{\infty}}),$
$\widetilde{B}=-\frac{8}{\beta_{0}^{2}}M\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu\|\psi\|_{\infty}})$
$\hskip 22.76228pt+\lambda T^{2}\varphi^{3}O(e^{4\mu\|\psi\|_{\infty}}),$